Replace the blank with an integer to make it a true statement. 
 _____ × (– 8) = –56 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us assume the missing integer be x,
Then,
= (x) × (-8) = -56
= x = (-56/-8)
= x =7
Let us substitute the value of x in the place of blank,
= (7) × (-8)
= -56 ... [∵(+ × - = -)]

Thanks